1.   Calcium phosphate, silicon dioxide, and coke may be heated together in an electric furnace to produce phosphorus, as shown in the following equation:

2 Ca3(PO4)2(s) + 6 SiO2(s) + 10 C(s) → 6 CaSiO3(s) + 10 CO(g) + P4(g)

In this reaction, what mass of calcium phosphate would be needed to make 100.0 g of phosphorus? What mass of calcium silicate would be formed as a by-product


1
Expert's answer
2021-05-10T01:50:39-0400

100 g= 45.3592 Kg of Phosphours P4


Molar mass of P4 = 123.90 gm/mol


Number of moles of P4 = 45359.2/123.90 = 366.09 moles


Moles of Ca3(PO4)2 required = 2 × moles of P4 = 2 × 366.09 = 732.19 moles


Molar mass of Ca3(PO4)2 = 310.18 gm/mol


Mass of Ca3(PO4)2 required = 732.19 × 310.18 gm = 227110.6942 gms = 227.110 Kg


Mass in Lbs = 500.69 Lbs


number of moles of CaSiO3 formed = 366.09 × 6 = 2196.54 moles


Molar mass of CaSiO3 = 124.33 gm/mol


Weight of CaSiO3 formed = 2196.54 × 124.33 = 273095.81 gms = 273.095 Kg = 602.07 pounds


number of moles of SiO2 required= 366.09 × 6 = 2196.54 moles


Molar mass of SiO2 = 60.08 gm/mol


Weight of SiO2 required = 2196.54 × 60.08 = 131968.12 gms = 131.968 Kg